OpenStudy (surana):

Determine the number of real solutions of X^2 +5x + 8 = 0

OpenStudy (surana):

The ^2 represents exponent 2 and the possible answers are A: 0 B: 1

OpenStudy (surana):

B is a double root, for notes.

OpenStudy (surana):

What I mean is that B's answer, 1, is a double root.

OpenStudy (jhannybean):

Use the discriminant, \(b^2-4ac\)

OpenStudy (surana):

How do I work that out?

OpenStudy (fibonaccichick666):

Google the fundamental theorem of Algebra, you should find it interesting :)

OpenStudy (anonymous):

b is the value attached to the x variable, a is the one attached to the x^2 variable and c is the value attached to no variable.

OpenStudy (surana):

So that means B would be 5x and C would be 8?

OpenStudy (fibonaccichick666):

^Value of the coefficients on those^^

OpenStudy (anonymous):

you could also use complete the square .

OpenStudy (anonymous):

No. b is 5 and c is 8.

OpenStudy (jhannybean):

\[b^2-4ac > 0 \implies 2~ \text{real solutions}\]\[b^2-4ac=0 \implies \text{1 real solution}\]\[b^2-4ac < 0 \implies \text{no real solutions}\]

OpenStudy (surana):

So I put 5 into B and 8 Into C?

OpenStudy (anonymous):

yes.

OpenStudy (jhannybean):

Know what your formula resembles. \[x^2+5x+8=0 \implies ax^2+bx+c=0\]

OpenStudy (fibonaccichick666):

And let me just emphasize where that all comes from: so, do you remember the quadratic formula? The discriminant is what is inside the square root

OpenStudy (surana):

So it should look like 25 - 4a 8?

OpenStudy (fibonaccichick666):

this is why we see if it is positive or negative

OpenStudy (surana):

But I think I need a number for a, right?

OpenStudy (jhannybean):

@Surana Can you identify what \[a=~?\]\[b=~?\]\[c=~?\]

OpenStudy (surana):

I think that A is supposed to be X^2.

OpenStudy (jhannybean):

Look at what I posted here and match up your formula with the quadratic. \[\color{red}{x^2+5x+8=0 \implies ax^2+bx+c=0}\]

OpenStudy (surana):

So I put 1 into the place of A, 5 into the place of B and 8 into the place of C?

OpenStudy (jhannybean):

So a b and c are called the coefficients of the variable terms. That is precisely what you are looking for to find the discriminant.

OpenStudy (jhannybean):

If nothing is in front of \(x^2\) you can infer that the variable coefficient =1.

OpenStudy (surana):

Then I need to calculate for A, B and C, first.

OpenStudy (surana):

I see what you mean now, so the A number is 1, since there was nothing explicitly mentioned there, meaning my answer is B?